# Attitudes Towards and Knowledge of Non-Consensual Synthetic Intimate Imagery in 10 Countries

### 26 Jan 2024 v1

@@ 5.2. Behaviors
The vast majority of respondents indicated no engagement in any type of NSII behaviors, with the most popular behavior being the consumption of celebrity deepfake pornography. It was even rarer for respondents to indicate perpetration themselves. This was similar for victimization, with few respondents indicating that they had experienced deepfake pornography. However, the victimization statistics in particular should be interpreted with caution. Even with traditional forms of image-based sexual abuse, many victims may be unaware of their victimhood. NSII may exacerbate this dynamic given that deepfake technology ”democratizes” victimization and image-based sexual abuse by enabling anyone with a computer to be a perpetrator, and anyone with imagery online to be targeted. Also, there are additional unique motivations for NSII that may result in less likelihood of victim notification; for instance, if the motivations for the creation or distribution of NSII are for sexual gratification purposes and/or for demonstrating or practicing digital skills in AI. Finally, as content may be distributed on dedicated sites or within closed communities on sites such as Discord or Reddit, victims of NSII may not know about or find out about their images.
@@
As a subset of image-based sexual abuse, it is useful to compare rates of NSII with previously reported statistics on image-based sexual abuse more broadly. There are several associated challenges, including sampling methodologies and definitions. For example, studies may define image-based sexual abuse behaviors differently, failing to include the non-consensual creation or filming of content, or threatening to share content, alongside the actual non-consensual sharing of content. Studies may also intentionally survey younger populations who may be more likely to have relevant experiences, as opposed to collecting population-level prevalence data. Notwithstanding those limitations, previous work has shown that image-based abuse prevalence is increasing over time, driven by changes in technology use, dating, sex, and privacy (Powell et al., 2020). Powell et al. (Powell et al., 2020) found on average, across Australia, New Zealand, and the UK, that  1/3 of survey respondents indicated at least one form of image-based sexual abuse, defined similarly to our study. An American-based study (Ruvalcaba and Eaton, 2020) focused exclusively on the non-consensual sharing of content, and found that 1/12 respondents indicated victimization, while 1/20 reported perpetration. Even taking the more expansive definition of image-based sexual abuse, our rates of perpetration and victimization using deepfake technology were quite low. Given the trends in traditional image-based sexual abuse (Powell et al., 2020), and the increasing ease of NSII creation, we expect these rates to increase over time.
5.3. Gender and NSII
@@
Awareness of NSII was higher among respondents who identified as men as compared to women. This is consistent with the literature on the ”digital divide” which suggests that women face disproportionate barriers to access and use to the internet compared to men. Even when women have access to the internet and technology, they may not have the time, resources, skills, or knowledge to leverage it (Acilar and Sæbø, 2023). Additionally, the attitudinal questions indicate a trend towards men diminishing the harms or ”badness” of NSII behaviors. This is consistent with previous findings on NSII (Fido et al., 2022), and with other findings around image-based sexual abuse more broadly (Attrill-Smith et al., 2021; Bothamley and Tully, 2017).
@@
Men were more likely to report both perpetration and victimization by NSII, across subtypes. Studies continue to report mixed findings regarding the role of gender in image-based sexual abuse. Powell et al. (Powell et al., 2020) reported higher rates of perpetration in men, and similar rates of victimization for men and women, but differential impacts. Our findings are somewhat inconsistent with a report which found the majority of NSII available online targeted women (Adjer et al., 2019). We argue that the differences in sampling methodology may help explain these inconsistencies. For instance, in the event of sexual extortion, the content may not be made available online. It is possible that the trend seen in gender may be somewhat explained by the motivation of the perpetrator, given previous research to support higher rates of ”sextortion” in men as compared to heterosexual women (Eaton et al., 2023), including for adolescent populations (Patchin and Hinduja, 2020). Another explanation for this discrepancy is that research suggests most deepfake videos online (an estimated 96%) are of women actors and musicians, who may be less likely to participate in survey research than non-public figures like our survey respondents (Adjer et al., 2019).
@@
While we asked broadly about motivations in the larger survey, we did not ask specifically ask about motivations for NSII perpetration or victimization. Motivations for NSII behaviors may range from personal sexual gratification, to use for extortion purposes, to demonstrating technical skill in the creation of content. A consideration for future research should be the use of NSII for extortion, particularly given anecdotal evidence of significant associated harms (e.g., suicide, suicidal ideation, or suicidal attempts) (Pu et al., 2021).
@@ 5.4. Limitations and Mitigation
Cross-country surveys are logistically difficult and require significant consideration in interpreting findings. Our study had to address language differences and non-response bias. The former were addressed, as detailed above, through cross-validation. The latter was a more significant concern, as others have noted that language used to discuss a topic such as image-based sexual abuse and pornography more broadly will be culturally dependent (Schoenebeck et al., 2023). We attempted to mitigate this by allowing respondents to select ”Prefer not to say” for any question that involved self-reporting on behavior, as well as periodically reminding respondents of that option throughout the survey. Nevertheless, it is possible that selection bias may have resulted in an underestimation of prevalence, with respondents opting out of the survey due to personal experiences or demographic characteristics. South Korea, for example, had a high percentage of bisexual respondents. This may indicate that our attempts at mitigating non-response bias were less successful in South Korea. We note that, despite explicit goals for representative sampling, the participant pool fell short, necessitating the need to weight results.
@@ 5.5. Design Considerations
In this study, we found that the majority of respondents across all countries were relatively unfamiliar with the concept of deepfake pornography. We also found that respondents, on average, found deepfake pornography behaviors to be deserving of punishment. Reported perpetration and victimization rates were low, with the most common behavior reported by respondents being the consumption of deepfake pornography. Finally, gender did play a part in both experiences and attitudes, with men deeming deepfake behaviors as less harmful or deserving of punishment as compared to women. This has implications for both policy and practice.
As relatively few respondents were aware of deepfake pornography, this signals first and foremost that more education is needed to both deter potential perpetration and/or the consumption of NSII. Education efforts should focus on the harms suffered by victim-survivors, as well as the potential consequences of perpetration. Given the gender dynamics in terms of perceived harmfulness, prioritization around education should focus on reaching boys and men.
Second, many scholars have focused on the need for legislation to criminalize these image-based abuse behaviors and provide clear redress for victims (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). In order for a law to have both general and specific deterrent effects, the consequences of breaking that law has to be known by the general public, and someone who does break the law has to be prosecuted and appropriately punished. Some of the trends, particularly the high rates of prevalence in both South Korea and Australia, indicate that laws are not a panacea to eliminate or even reduce this behavior. For instance, Australia criminalizes the distribution of non-consensual intimate imagery, including NSII (Australia, 2021), while South Korea has laws allowing for up to 3 years imprisonment for merely viewing non-consensual intimate imagery (Center, 2019; Maeng and Lee, 2022). Given that this behavior primarily happens online, that victim-survivors may not be aware of the material, or if they become aware, may be reluctant to report it, it can be argued that existing legislation provides weak deterrence since very few people have been prosecuted for NSII in jurisdictions that do have such laws in place. As we found in this study, there was also low awareness of the concept of deepfake pornography, and one may reasonably infer low awareness of related laws. While developing legislation around an issue helps create societal norms and signals the unacceptability of a behavior, it is crucial to concede the limitations of passing regulations in enacting true societal change. This also allows us to develop parallel efforts, such as technical solutions, that may mitigate these harms by reducing the creation and dissemination of NSII.
Third, there are a number of points of potential intervention for technical solutions. First, discovery tools that allow individuals to search the web for sexual material via facial recognition technology may be empowering, although there are safety, privacy, and security concerns that need to be taken into account (e.g., building in protections such that individuals can only search for content of their own faces). Second, while many videos are posted on dedicated deepfake websites, or make no claims as to veracity, the use of deepfakes for sexual extortion indicate that technical methods of annotating content, such as watermarks, may help mitigate NSII harms, although not fully (Harris, 2021). Third, platforms can make choices around what they allow and disallow in their policies and community standards (legislation can assist in justifying the choices made by platforms). Google Search (goo, 2023), Reddit (Hollister, 2023), and PornHub(Cole, 2018) have already published dedicated NSII policies. Fourth, various stakeholders have worked to help create solutions for victims, primarily focusing on triaging after victimization has occurred. It will be necessary for these help-seeking resources (whether helplines, online forms, or chatbots(Falduti and Tessaris, 2022; Maeng and Lee, 2022)) to be aware of this new type of image-based sexual abuse and up-to-date with the types of redress that victims may seek (Aliman et al., 2021). Fifth, where tools do have a payment or sign up flow, users could be required as part of the flow to ingest and acknowledge information about the harms and potential consequences of NSII creation and dissemination. Finally, platforms have demonstrated an ability to detect nude content or, in some cases, non-consensual nude or sexual content. If creator tools could use those tools to block creation, or at the very least serve a warning interstitial about the harms associated with NSII, it may act as a further deterrent for uploading this content.
@@
As practitioners and designers think about future considerations for intervention, it may be helpful to break this issue down into relevant behaviors as seen in Figure 7 and consider intervention options (which should be considered complementary, as opposed to solely sufficient) at each point in the process. The increased ease of access to deepfake technologies was the catalyst for this project, and represents the first place at which interventions could be made. Are there ways, for example, to limit these tools’ abilities to create nude or intimate imagery? If creators get to the point where barriers have been hacked, are there warning messages that can either highlight associated harms to targets, or potential punishments associated with creation/distribution of such content? After creation, are there ways to build in indicators of provenance (e.g., watermarking) that signal the synthetic nature of the content? Finally, assuming most impediments can be circumvented by a motivated creator, empathetic and multi-modal resources for victims could include both tooling to discover where relevant content is published, and help for seeking take downs and recourse more broadly. These are primarily design interventions, but overarching all of these would be the societal interventions of educational programming, and relevant legislation making clear potential punishments associated with each type of behavior (extortion, posting/distribution, etc); paired with prosecutorial actions that indicate a willingness to hold perpetrators to account.
@@ 5.6. Conclusion
In this study, we used a survey to explore general attitudes toward NSII, as well as NSII behaviors and experiences in ten different countries around the world. We found relatively low rates of awareness across countries, but respondent attitudes (as seen in the questions about whether victims should get upset and criminalization, and in the qualitative responses) were concern and condemnation of image-based sexual abuse behaviors, including NSII. The most popular reported behavior was the consumption of celebrity NSII, while personal victimization and perpetration rates were quite low. Men tended to find consuming NSII less objectionable than women, which was also seen in the higher rates of self-reported consumption behavior. Finally, men were more likely to report both more perpetration and more victimization by NSII, a trend that future research should explore as we may expect prevalence to increase over time. Laws, policies, and support services can help to mitigate the harms of NSII; however, more resources need to be directed toward detecting and blocking the creation of content prior to dissemination, as well as preventing the abuse from happening in the first place through educational programs and initiatives.
@@
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.
@@
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.

### 13 Feb 2024 v2

@@ 5.2. Behaviors
The vast majority of respondents indicated no engagement in any type of AI-IBSA behaviors, with the most popular behavior being the consumption of celebrity deepfake pornography. It was even rarer for respondents to indicate perpetration themselves. This was similar for victimization, with few respondents indicating that they had been victimized by deepfake pornography. However, the victimization statistics in particular should be interpreted with caution. Even with traditional forms of image-based sexual abuse, many victims may be unaware of their victimhood. AI-IBSA may exacerbate this dynamic given that deepfake technology “democratizes” victimization and image-based sexual abuse by enabling anyone with a computer to be a perpetrator, and anyone with imagery online to be targeted. Also, there are additional unique motivations for AI-IBSA that may result in less likelihood of victim notification; for instance, if the motivations for the creation or distribution of AI-IBSA are for sexual gratification purposes and/or for demonstrating or practicing digital skills in AI. Finally, as content may be distributed on dedicated sites or within closed communities on sites such as Discord or Reddit, victims of AI-IBSA may not know about or find out about their images.
@@
As a subset of image-based sexual abuse, it is useful to compare rates of AI-IBSA with previously reported statistics on image-based sexual abuse more broadly. There are several associated challenges, including sampling methodologies and definitions. For example, studies may define image-based sexual abuse behaviors differently, failing to include the non-consensual creation or filming of content, or threatening to share content, alongside the actual non-consensual sharing of content. Studies may also intentionally survey younger populations who may be more likely to have relevant experiences, as opposed to collecting population-level prevalence data. Notwithstanding those limitations, previous work has shown that image-based abuse prevalence is increasing over time, driven by changes in technology use, dating, sex, and privacy (Powell et al., 2020). Powell et al. (Powell et al., 2020) found on average, across Australia, New Zealand, and the UK, that  1/3 of survey respondents indicated at least one experience of one or more forms of image-based sexual abuse, defined similarly to our study. An American-based study (Ruvalcaba and Eaton, 2020) focused exclusively on the non-consensual sharing of content, and found that 1/12 respondents indicated victimization, while 1/20 reported perpetration. Even taking the more expansive definition of image-based sexual abuse, our rates of perpetration and victimization using deepfake technology were quite low. Given the trends in traditional image-based sexual abuse (Powell et al., 2020), and the increasing ease of NSII creation, we expect these rates to increase over time.
5.3. Gender and AI-IBSA
@@
Awareness of deepfakes was higher among respondents who identified as men as compared to women. This is consistent with the literature on the “digital divide” which suggests that women face disproportionate barriers to access and use to the internet compared to men. Even when women have access to the internet and technology, they may not have the time, resources, skills, or knowledge to leverage it (Acilar and Sæbø, 2023). Additionally, the attitudinal questions indicate a trend towards men diminishing the harms or “badness” of deepfake pornography behaviors. This is consistent with previous findings on deepfake pornography (Fido et al., 2022), and with other findings around image-based sexual abuse more broadly (Attrill-Smith et al., 2022; Bothamley and Tully, 2017).
@@
Men were more likely to report both perpetration and victimization by deepfake pornography, across subtypes. Studies continue to report mixed findings regarding the role of gender in image-based sexual abuse. Powell et al. (Powell et al., 2020) reported higher rates of perpetration in men, and similar rates of victimization for men and women, but differential impacts. Our findings are somewhat inconsistent with a report which found the majority of AI-IBSA available online targeted women (Adjer et al., 2019). We argue that the differences in sampling methodology may help explain these inconsistencies. For instance, in the event of sexual extortion, the content may not be made available online. It is possible that the trend seen in gender may be somewhat explained by the motivation of the perpetrator, given previous research to support higher rates of “sextortion” in men as compared to heterosexual women (Eaton et al., 2023), including for adolescent populations (Patchin and Hinduja, 2020). Another explanation for this discrepancy is that research suggests most deepfake videos online (an estimated 96%) are sexualized images of women actors and musicians, who may be less likely to participate in survey research than non-public figures like our survey respondents (Adjer et al., 2019).
@@
While we asked broadly about motivations in the larger survey, we did not ask specifically ask about motivations for deepfake pornography perpetration or victimization. Motivations for such behaviors may range from personal sexual gratification, to use for extortion purposes, to demonstrating technical skill in the creation of content. A consideration for future research should be the use of AI-IBSA (as well as other forms of NSII not involving AI) for extortion, particularly given anecdotal evidence of significant associated harms (e.g., suicide, suicidal ideation, or suicidal attempts) (Pu et al., 2021).
@@ 5.4. Limitations and Mitigation
Cross-country surveys are logistically difficult and require significant consideration in interpreting findings. Our study had to address language differences and non-response bias. The former were addressed, as detailed above, through cross-validation. The latter was a more significant concern, as others have noted that language used to discuss a topic such as image-based sexual abuse and pornography more broadly will be culturally dependent (Schoenebeck et al., 2023). We attempted to mitigate this by allowing respondents to select “Prefer not to say” for any question that involved self-reporting on behavior, as well as periodically reminding respondents of that option throughout the survey. Nevertheless, it is possible that selection bias may have resulted in an underestimation of prevalence, with respondents opting out of the survey due to personal experiences or demographic characteristics. South Korea, for example, had a high percentage of bisexual respondents. This may indicate that our attempts at mitigating non-response bias were less successful in South Korea. In addition, in this study, we evaluated prevalence based on respondents who indicated victimization with deepfake pornography specifically. Thus, we may have underestimated prevalence because we did not include respondents who were unsure, or who incorrectly indicated that their content was not deepfaked. Finally, we note that, despite explicit goals for representative sampling, the participant pool fell short, necessitating the need to weight results.
@@ 5.5. Design Considerations
In this study, we found that the majority of respondents across all countries were relatively unfamiliar with the concept of deepfake pornography. We also found that respondents, on average, found deepfake pornography behaviors to be deserving of punishment. Reported perpetration and victimization rates were low, with the most common behavior reported by respondents being the consumption of deepfake pornography. Finally, gender did play a part in both experiences and attitudes, with men deeming deepfake pornography behaviors as less harmful or deserving of punishment as compared to women. This has implications for both policy and practice.
As relatively few respondents were aware of deepfake pornography, this signals first and foremost that more education is needed to both deter potential perpetration and the consumption of AI-IBSA. Education efforts should focus on the harms suffered by victim-survivors, as well as the potential consequences of perpetration. Given the gender dynamics in terms of perceived harmfulness, prioritization around education should focus on reaching boys and men.
@@
Second, many scholars have focused on the need for legislation to criminalize these image-based abuse behaviors and provide clear redress for victims (Chesney and Citron, 2019; Gieseke, 2020; Delfino, 2019; Mania, 2020; Solichah et al., 2023). In order for a law to have both general and specific deterrent effects, the consequences of breaking that law has to be known by the general public, and someone who does break the law has to be prosecuted and appropriately punished. Some of the trends—particularly the high rates of prevalence in both South Korea and Australia—indicate that laws are not a panacea to eliminate or even reduce this behavior. For instance, some Australian states and territories specifically criminalize the distribution of NSII (Australia, 2021), while South Korea has laws allowing for up to 3 years imprisonment for merely viewing non-consensual intimate imagery (Center, 2019; Maeng and Lee, 2022). Given that this behavior primarily happens online, that victim-survivors may not be aware of the material, or if they become aware, may be reluctant to report it, it can be argued that existing legislation provides weak deterrence since very few people have been prosecuted for NSII in jurisdictions that do have such laws in place. As we found in this study, there was also low awareness of the concept of deepfake pornography, and one may reasonably infer low awareness of related laws. While developing legislation around an issue helps create societal norms and signals the unacceptability of a behavior, it is crucial to concede the limitations of passing regulations in enacting true societal change. The limitations of law underscores the importance of developing parallel efforts, such as technical solutions, that may mitigate these harms by reducing the creation and dissemination of NSII.
Third, there are a number of points of potential intervention for technical solutions. First, discovery tools that allow individuals to search the web for sexual material via facial recognition technology may be empowering, although there are safety, privacy, and security concerns that need to be taken into account (e.g., building in protections such that individuals can only search for content of their own faces). Second, while many videos are posted on dedicated deepfake websites, or make no claims as to veracity, the use of deepfakes for sexual extortion indicate that technical methods of annotating content, such as watermarks, may help mitigate NSII harms (Harris, 2021). Third, platforms can make choices around what they allow and disallow in their policies and community standards (legislation can assist in justifying the choices made by platforms). Google Search (Google, 2023), Reddit (Hollister, 2023), and PornHub(Cole, 2018) have already published dedicated NSII policies. Fourth, various stakeholders have worked to help create solutions for victims, primarily focusing on triaging after victimization has occurred. It will be necessary for these help-seeking resources (whether helplines, online forms, or chatbots (Falduti and Tessaris, 2022; Maeng and Lee, 2022)) to be aware of this new type of image-based sexual abuse and up-to-date with the types of redress that victims may seek (Aliman et al., 2021). Fifth, where tools do have a payment or sign up flow, users could be required as part of the flow to ingest and acknowledge information about the harms and potential consequences of NSII creation and dissemination. Finally, platforms have demonstrated an ability to detect nude content or, in some cases, non-consensual nude or sexual content. If creator tools could use those tools to block creation, or at the very least serve a warning interstitial about the harms associated with NSII (including AI-IBSA), it may act as a further deterrent for uploading this content.
@@
As practitioners and designers think about future considerations for intervention, it may be helpful to break this issue down into relevant behaviors as seen in Figure 7 and consider intervention options (which should be considered complementary, as opposed to solely sufficient) at each point in the process. The increased ease of access to NSII creation technologies was the catalyst for this project, and represents the first place at which interventions could be made. Are there ways, for example, to limit these tools’ abilities to create nude or intimate imagery? If creators get to the point where barriers have been hacked, are there warning messages that can either highlight associated harms to targets, or potential punishments associated with the creation or distribution of such content? After creation, are there ways to build in indicators of provenance (e.g., watermarking) that signal the synthetic nature of the content? Finally, assuming most impediments can be circumvented by a motivated creator, empathetic and multi-modal resources for victims could include both tooling to discover where relevant content is published, and help for seeking take downs and recourse more broadly. These are primarily design interventions, but overarching all of these would be the societal interventions of educational programming, and relevant legislation making clear potential punishments associated with each type of behavior (extortion, posting/distribution, etc), paired with prosecutorial actions that indicate a willingness to hold perpetrators to account.
@@ 5.6. Conclusion
In this study, we used a survey to explore general attitudes toward AI-IBSA (or “deepfake pornography”), as well as AI-IBSA behaviors and experiences in 10 different countries around the world. We found relatively low rates of awareness across countries, but respondents expressed concern and condemnation of image-based sexual abuse behaviors, including deepfake pornography (as seen in the questions about whether victims should get upset and criminalization, and in the qualitative responses). The most popular reported behavior was the consumption of celebrity deepfake pornography, while personal victimization and perpetration rates were quite low. Men tended to find consuming deepfake pornography less objectionable than women, which was also seen in the higher rates of self-reported consumption behavior. Finally, men were more likely to report both more perpetration and more victimization by deepfake pornography, a trend that future research should explore as we may expect prevalence to increase over time. Laws, policies, and support services can help to mitigate the harms of NSII; however, more resources need to be directed toward detecting and blocking the creation of content prior to dissemination, as well as preventing the abuse from happening in the first place through educational programs and initiatives.
@@
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.
@@
Figure A3. Heatmap of Median Criminalization Attitudes by Gender. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Numbers shown are medians.